OpenStudy (anonymous):

Simplify 4 plus 2 over x y all over 2

OpenStudy (anonymous):

looks like this 

OpenStudy (ash2326):

We have \[\huge \frac{4+\frac{2}{xy}}{2}\] First step is to multiply numerator and denominator by xy \[\huge \frac{4\times xy+\frac{2}{xy}\times xy}{2\times xy}\] What would you get now?

OpenStudy (anonymous):